OpenStudy (anonymous):

. A sphere is inscribed in a cube. Find the lateral area of the cube if the area of the sphere is 784π.

OpenStudy (anonymous):

Did you know the formula for the area of the sphere ?

OpenStudy (anonymous):

4 pi r squared

OpenStudy (anonymous):

Can u find the radius r, from te give question?

OpenStudy (mayankdevnani):

do you know what is the lateral surface area of cube

OpenStudy (mayankdevnani):

4b^2=lateral surface area of cube

OpenStudy (anonymous):

the radius is 4.998732447

OpenStudy (mayankdevnani):

like r=4.9

OpenStudy (anonymous):

right i think so.

OpenStudy (anonymous):

dis gon b gud

OpenStudy (anonymous):

wat do u think is the answer? pls help me that`s our take home finals tnx.

OpenStudy (anonymous):

\[A_{sphere}=4{\pi}r^{2}\]\[784\pi=4{\pi}r^{2}\]\[r^{2}=\frac{ 784 }{ 4 }\]\[r=14\]Since r is half of the side of the cube, the length of the side of the cube (b) is equal to 2r . Thus \[A_{l}=5b^{2}=5(2r)^{2}=5(2\times14)^{2}=3920\]

OpenStudy (anonymous):

so that is the lateral area of the cube?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

@mayankdevnani can you confirm?

OpenStudy (anonymous):

thank you, do you think you can help me with the other one? if not its ok

OpenStudy (mayankdevnani):

what? @Algebraic!

OpenStudy (anonymous):

. Suppose the lateral area of a right circular cone and the area of a sphere equal 64π in2. If the radius of the sphere and the radius of the base of the cone are equal, what is the volume of the cone? (10 points)

OpenStudy (mayankdevnani):

@Zekarias plz tell me what is lateral surface area of cube

OpenStudy (anonymous):

L.S.A of a cube = 4l^2

OpenStudy (anonymous):

@Zekarias not 5l^2

OpenStudy (mayankdevnani):

yaa that's what i called this question

OpenStudy (anonymous):

LSA of a regular object is Perimeter of base times height.

OpenStudy (anonymous):

sorry you are right. It 4b^{2}

OpenStudy (anonymous):

so what is the L S A of the cube?

OpenStudy (anonymous):

sauravshakya so how is the solution??

OpenStudy (anonymous):

\[4(2r)^{2}\]

OpenStudy (anonymous):

@Zekarias would like to do that.

OpenStudy (mayankdevnani):

4(2r^2)=4(2*14^2)=1568

OpenStudy (mayankdevnani):

ok @dylan18

OpenStudy (anonymous):

ok @Zekarias @ sauravshakya

OpenStudy (anonymous):

@mayankdevnani that is the soulution?

OpenStudy (anonymous):

NOPE @mayankdevnani

OpenStudy (mayankdevnani):

yaa

OpenStudy (anonymous):

LSA of a cube =4 l^2 = 4*(2*14)^2 = 3136

OpenStudy (anonymous):

@Zekarias i think he can do that

OpenStudy (anonymous):

square units

OpenStudy (anonymous):

i mean you @Zekarias

OpenStudy (mayankdevnani):

yaa sry

OpenStudy (mayankdevnani):

thnx.. @sauravshakya

OpenStudy (anonymous):

ah... the unit is square units.. is that the answer?

OpenStudy (mayankdevnani):

yaa

OpenStudy (anonymous):

ahhh., thank you very much

OpenStudy (anonymous):

thank you very much @Zekarias ,@sauravshakya ,@mayankdevnani you think you can help me last one.?

OpenStudy (mayankdevnani):

welcome

OpenStudy (anonymous):

6. Suppose the lateral area of a right circular cone and the area of a sphere equal 64π in2. If the radius of the sphere and the radius of the base of the cone are equal, what is the volume of the cone? (10 points) tnx

OpenStudy (anonymous):

you think that is enough @Zekarias ?

OpenStudy (anonymous):

What is the question now?

OpenStudy (anonymous):

6. Suppose the lateral area of a right circular cone and the area of a sphere equal 64π in2. If the radius of the sphere and the radius of the base of the cone are equal, what is the volume of the cone? (10 points)

OpenStudy (anonymous):

Ok to find the volume of a cone you need to have the height (h) and radius of the base (r). To do that u need to find the slant height (l). Did u know the formula for the lateral surface area of a cone?

OpenStudy (anonymous):

i think the LSA is in the question.

OpenStudy (anonymous):

It the voume what u asked Any how .....here is what I asked u befor\[A_{l..cone}={\pi}rl...and .....A_{sphere}=4{\pi}r^{2}\]

OpenStudy (anonymous):

\[A_{l..cone}={\pi}rl=A_{sphere}=4{\pi}r^{2}\]since bothe are equal

OpenStudy (anonymous):

π × r × l, like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then how should we used that?

OpenStudy (anonymous):

Now our goal is to find l first. ok

OpenStudy (anonymous):

yes.. i think we can used base*height?

OpenStudy (anonymous):

\[A_{sphere}=4{\pi}r^{2}=64\pi\]\[r^{2}=16\]\[r=4\]and \[A_{l..cone}={\pi}rl=64\pi......but ..r=4\]\[4l=64\]\[l=16\]

OpenStudy (anonymous):

 Now can u find h

OpenStudy (anonymous):

SO NOW WE can substitute to find the height

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

U get this \[h=4\sqrt{15}\]

OpenStudy (anonymous):

15.49

OpenStudy (anonymous):

Now the volume of the cone is given by\[V=\frac{ 1 }{ 3 }{\pi}r^{2}h\]

OpenStudy (anonymous):

can u finish it?

OpenStudy (anonymous):

(1/3) * pi * r2 * h - or - 3.14159 *r2 * l * 4/3

OpenStudy (anonymous):

correct i out now GOOD BY

OpenStudy (anonymous):

thanks